OpenStudy (anonymous):

the line that is described by the equation y=-4x+1 contains which of the following A) (0,1) b) (-3,1) C) (1,3) D) (-2,-3)

OpenStudy (anonymous):

You can check fairly easily. Take the first number in the parenthesis. That is what you plug in for x. Then you just see what y is. If it's the same as the second number in the parenthesis, that point is on the line! As an example, if you had (9,7), you would say y=-4*9+1 y=-36+1 y=-35 And since 7=/=-35, you know that (9,7) is not on the line.